OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 7 - 9x - (x + 12) <= 25?

myininaya (myininaya):

-9x-x-12<=25 -10x-12<=25 -10x<=37 x>=-37/10 what ever solution satisfies x>=-37/10

OpenStudy (anonymous):

Thats not on option

OpenStudy (anonymous):

?

myininaya (myininaya):

?

OpenStudy (anonymous):

Thats not one of the choices I was givin

myininaya (myininaya):

the question says which of the ways is not a way to represent the solution there are many ways to not represent the solution incorrectly

myininaya (myininaya):

i mean there are many ways to represent the solution incorrectly*

OpenStudy (anonymous):

x >= -3 x <= -3 -3 <= x A number line with a closed circle on negative 3 and shading to the right

myininaya (myininaya):

-9x-x-12<=25 this is the inequality i solved -10x-12<=25 -10x<=37 x>=-37/10